If, in a sample of n = 16 selected from a normal population, x bar = 57, and s = 8, what are the critical values of t if the level of significance, is .01, the null hypothesis h0 is mu = 50, and the alternative hypothesis H1 is mu is not equal to 50.

The critical values of t are +/- ___ , ____.
Answer #1


Degree of freedom = n-1 = 16-1 = 15

t-critical is calculated by looking at t-distribution tables for degrees of freedom 15 and significance level 0.01

This gives t-critical = +/-2.9467 (two tailed test)
Answer #2

SOLUTION :


DF = 16 - 1 = 15 


For 0.01 level of significance, critical t-values from t-distribution at DF of 15 is : 

- 2.947 and + 2.947  (ANSWER).

answered by: Tulsiram Garg